OpenStudy (anonymous):

1.Use order of operations and the mathematical properties of numbers to simplify these numbers. 8 - [(8 - 5) x 2] = Type the correct answer. 2.Use order of operations and the mathematical properties of numbers to simplify these numbers. 6 x 7 - 7 x 3 = Type the correct answer.

OpenStudy (anonymous):

Well, the order of operations is an acronym called PEDMAS... dO U KNOW THIS?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

PEMDAS

OpenStudy (anonymous):

MY MISTAKE LOL

OpenStudy (anonymous):

Parenthesis, exponents, multipication, division, addition, subtraction

OpenStudy (anx):

Use Pemdas P (Parenthesis) E (Exponent) M (Mulitply) D (Divide) A (Add) S (subtrat) From LEFT TO RIGHT so for 6 x 7 - 7 x3 (6x7) = 42 - (7 x3 ) = 21 42 - 21= 21

OpenStudy (anonymous):

Just like @anx said!

OpenStudy (anx):

For the first one 8 - [(8-5) x2 ] Bracket first Then parenthesis so 8 - [(3)x2) so 8 - (6) = x

OpenStudy (anx):

Should be x = 2, sorry.

OpenStudy (anonymous):

thank u

OpenStudy (anx):

Anytime.